OpenStudy (anonymous):

solve: 2sin(3x)=1 on [-pi,pi] thx

OpenStudy (anonymous):

/2 so you get sin(3x)=1/2

OpenStudy (anonymous):

now you can think of this as sin x = 1/2 which will be pi/6 but you have 3 theta so pi/6*3

OpenStudy (anonymous):

sin(3x)=1/2 since sin(x)=1/2 at pi/6 and 5pi/6 3x=pi/18 and 3x=5pi/18

OpenStudy (anonymous):

so your answers will be pi/18 and 5pi/18